OpenStudy (anonymous):

(x+1)(x^2+2x-4) how to multiply this polynomial?

OpenStudy (anonymous):

We distribute the x, then the 1: \((x+1)(x^2+2x-4) =x^3+2x^2-4x+x^2+2x-4=x^3+3x^2-2x-4 \)

jimthompson5910 (jim_thompson5910):

You can think of (x+1)(x^2+2x-4) as x(x^2+2x-4) + 1(x^2+2x-4)

jimthompson5910 (jim_thompson5910):

Distribute that and combine like terms to simplify

OpenStudy (anonymous):

yeah I got it thanx. my answer is correct

jimthompson5910 (jim_thompson5910):

yw